# Offline Clustering of Linear Bandits: Unlocking the Power of Clusters in Data-Limited Environments

### v1

@@ 4.3 Strategies for Choosing $\hat{\gamma}$
As demonstrated in Theorem 4.2, the performance of Algorithm 1 critically depends on the choice of $\hat{\gamma}$. A carefully chosen $\hat{\gamma}$ balances homogeneous and heterogeneous errors, directly affecting the algorithm’s suboptimality. In this subsection, we outline strategies for selecting $\hat{\gamma}$ under different scenarios. Figure 1 shows the influence of different choices for $\hat{\gamma}$.
@@
Case 1: Exact Information of $\gamma$. We set $\hat{\gamma}=\gamma$ when the exact value of $\gamma$ is known. This ensures $\varepsilon=0$ in Theorem 4.2, yielding only users from the same true cluster as $u_{\textnormal{test}}$ are included, i.e. $\eta_{\mathcal{W}_{\gamma}(u_{\textnormal{test}})}=0$. The resulting suboptimality is formalized in Corollary 4.5.
@@
With the same notations as Theorem 4.2, if $\hat{\gamma}=\gamma$ in Algorithm 1, the suboptimality gap satisfies:
[…]
@@
Corollary 4.5 implies that Algorithm 1 achieves suboptimality only with homogeneous error. This is because no users from outside $u_{\textnormal{test}}$’s cluster are included, eliminating heterogeneous error entirely. Furthermore, Algorithm 1 only connects homogeneous users based on similar preference vectors, avoiding the biased inclusion of heterogeneous users.
@@
$\Gamma(u,v)$ $=\left\|\hat{\bm{\theta}}_{u}-\hat{\bm{\theta}}_{v}\right\|_{2}-\alpha(\text{CI}_{u}+\text{CI}_{v}),\ M(u)=\{v\in\mathcal{V}\setminus\{u\}\mid\Gamma(u,v)>0\}$ (4)
where $\text{CI}_{u}$ denotes the confidence bound for user $u$’s estimated preference vector. Since $\Gamma(u,v)\leq\left\|\bm{\theta}_{u}-\bm{\theta}_{v}\right\|_{2}$ for $\alpha\geq 1$, it serves as a lower bound on the true difference between users’ preference vectors. If $\Gamma(u_{1},u_{2})\leq 0$, it suggests either the users belong to the same cluster or the available data is insufficient to distinguish them. Therefore, the set $M(u)$ includes users that are likely to be from different clusters than $u$, while those outside $M(u)$ either share the same preferences or cannot be confidently separated due to limited data.
Below, we propose two policies for selecting $\hat{\gamma}$ based on the users in $M(u_{\textnormal{test}})$. Note that in cases where $M(u_{\textnormal{test}})$ is empty, it is natural to choose $\hat{\gamma}=0$, as no sufficient data exists to identify any property of clusters. By selecting $\hat{\gamma}=0$, Algorithm 1 degenerates to LinUCB (Abbasi-Yadkori et al., 2011), using only the samples from $\mathcal{D}_{u_{\textnormal{test}}}$.
Underestimation Policy. The underestimation policy for choosing $\hat{\gamma}$ is defined as:
$\textstyle\hat{\gamma}=\mathbb{I}\{M(u_{\textnormal{test}})\neq\emptyset\}\cdot\min_{v\in M(u_{\textnormal{test}})}\Gamma(u_{\textnormal{test}},v).$ (5)
The intuition behind this approach stems from the property that $\Gamma(u,v)\leq\left\|\bm{\theta}_{u}-\bm{\theta}_{v}\right\|_{2}$. Since all users in $M(u_{\textnormal{test}})$ are from clusters different from that of $u_{\textnormal{test}}$, the underestimation ensures that $\hat{\gamma}$ is sufficiently small, thus avoiding the inclusion of too many biased samples in decision-making. This reduces the risk of introducing a high heterogeneous error, which explains why the minimum value of $\Gamma(u_{\textnormal{test}},v)$ is chosen from all users in $M(u_{\textnormal{test}})$. To formalize this, we present Lemma 4.7, which demonstrates how the underestimation policy can effectively restrict the size of $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$.
Lemma 4.7 (Effect of Underestimation).
For $\hat{\gamma}$ chosen according to Equation 5 and $\alpha\geq 1$, any user $v$ within the set for estimating heterogeneous neighbors $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$ in Theorem 4.2 also satisfies that $K_{u_{\text{test}}}+K_{v}\geq\alpha_{r}\left\|\bm{\theta}_{u_{\text{test}}}-\bm{\theta}_{v}\right\|_{2}$ for $\alpha_{r}$ defined in Theorem 4.2.
Lemma 4.7 indicates that choosing $\hat{\gamma}$ as specified in Equation 5 effectively limits the size of $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$. The condition $K_{u_{\textnormal{test}}}+K_{v}\geq\alpha_{r}\|\bm{\theta}_{u_{\textnormal{test}}}-\bm{\theta}_{v}\|_{2}$ ensures that only heterogeneous users without much data are included in $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$, further reducing the impact of biased samples. However, while this policy is effective at restricting $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$, the small value of $\hat{\gamma}$ also limits the size of $\mathcal{R}_{\hat{\gamma}}(u_{\textnormal{test}})$, the set of users within the same cluster as $u_{\textnormal{test}}$. This reduction in sample size may affect the quality of estimation and increase the homogeneous error. Therefore, the underestimation policy is most suitable in scenarios where avoiding biased samples and minimizing heterogeneous error are critical, even at the expense of reducing the number of available samples within the same cluster. For example, in medical treatment scenarios, introducing heterogeneous error could pose significant risks to patients, making it critical to prioritize avoiding such errors over utilizing more available samples.
Overestimation Policy. In contrast to underestimation, the overestimation policy uses an upper-bound estimate to choose $\hat{\gamma}$. The policy is defined as:
@@
$\textstyle\hat{\gamma}=\mathbb{I}\{M(u_{\textnormal{test}})\neq\emptyset\}\cdot\min_{v\in M(u_{\textnormal{test}})}\tilde{\Gamma}(u_{\textnormal{test}},v),$ (6)
where $\tilde{\Gamma}(u_{\textnormal{test}},v)=\left\|\hat{\bm{\theta}}_{u_{\textnormal{test}}}-\hat{\bm{\theta}}_{v}\right\|_{2}+\alpha(\text{CI}_{u_{\textnormal{test}}}+\text{CI}_{v})$ is an upper-bound estimate of the gap between the test user $u_{\textnormal{test}}$ and another user $v$.
Lemma 4.8 (Effect of Overestimation).
For $\hat{\gamma}$ chosen as in Equation 6, if $M(u_{\text{test}})\neq\emptyset$, $\hat{\gamma}\geq\gamma$.
By choosing $\hat{\gamma}\geq\gamma$ according to Equation 6, Algorithm 1 ensures that samples from homogeneous users with sufficient data are included in $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})$. This significantly reduces homogeneous error by leveraging a larger dataset. However, as $\hat{\gamma}\geq\gamma$, this policy may also introduce some biased users into $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$, increasing the heterogeneous error. The trade-off here is between reducing homogeneous error and potentially increasing heterogeneous error. Therefore, the overestimation policy is particularly effective in scenarios where combining a large amount of data from diverse users is critical, e.g., for high-dimensional preference vectors, and the goal is to minimize homogeneous error rather than strictly controlling heterogeneous error.
5 Algorithm for Sufficient Dataset: Offline - Clustering of Bandits
@@
When $\gamma$ is unknown, Algorithm 1 (Off-C2LUB) faces a trade-off: setting $\hat{\gamma}$ too small increases homogeneous error, while setting it too large introduces heterogeneous error, as discussed in Section 4.3. To address this, we propose a second algorithm, Offline Clustering of Bandits (Off-CLUB), that avoids explicitly estimating $\gamma$ and achieves strong performance when the dataset is sufficiently large.
@@ 5.1 Algorithm 2: Off-CLUB
Off-CLUB shares the same two-phase structure as Off-C2LUB, consisting of a Cluster Phase and a Decision Phase. Due to space constraints, full details and pseudo-code (Algorithm 2) are deferred to Appendix D. The main difference lies in graph construction: Off-CLUB starts with a complete graph and removes edges between users with dissimilar preferences, whereas Off-C2LUB begins with an empty graph and adds edges only when sufficient similarity is detected.
@@
$\textstyle\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}>\alpha\left(\text{CI}_{u_{1}}+\text{CI}_{u_{2}}\right),$ (7)
indicating that users $u_{1}$ and $u_{2}$ likely belong to different clusters. After edge removal, the resulting graph $\tilde{\mathcal{G}}$ correctly clusters users with sharing preference vectors, provided that all users have sufficient samples as discussed in Section 5.2. Finally, the subsequent steps, including data aggregation, computation of statistics, and the Decision Phase, are largely similar to those in Off-C2LUB.
@@ 5.2 Upper Bound for Off-CLUB under Sufficient Data
The effectiveness of Off-CLUB depends on correctly removing edges between users in different clusters, which requires sufficiently small confidence bounds $\text{CI}_{u}$. This condition is typically met when each user has a large number of samples. Conversely, users with limited samples may remain incorrectly connected, causing users from different clusters to be grouped together. Thus, to formalize the conditions for Off-CLUB’s success, we introduce the sufficient data assumption:
@@
The dataset $\mathcal{D}_{u}$ is called $\delta$-sufficient for user $u$ if $N_{u}\geq\max\left\{\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8dU}{\lambda_{a}^{2}\delta}\right),\,\frac{512d}{\gamma^{2}\lambda_{a}}\log\left(\frac{2U}{\delta}\right)\right\}$ for some $\delta\in(0,1)$. Furthermore, we assume that the full dataset $\mathcal{D}=\{\mathcal{D}_{u}\}_{u=1}^{U}$ is completely $\delta$-sufficient, i.e., each user’s dataset $\mathcal{D}_{u}$ is $\delta$-sufficient.
This assumption ensures that each user has enough samples to accurately estimate preference vectors, enabling Off-CLUB to effectively cluster users. Theorem 5.2 (proof in Appendix A.4) establishes the performance guarantees of Off-CLUB under this assumption.
@@
For any $\delta\in(0,\frac{1}{2})$, assume $\mathcal{D}$ is completely $\delta$-sufficient. For a given input $\alpha=1$ and $\lambda>0$, with probability at least $1-2\delta$, the suboptimality gap satisfies
$\textnormal{SubOpt}(u_{\textnormal{test}},\textnormal{Off-CLUB},A_{\textnormal{test}})\leq\tilde{O}\left(\sqrt{\frac{d}{N^{\mathcal{V}(j_{u_{\textnormal{test}}})}}}\right)$
where $N^{\mathcal{V}(j_{u_{\textnormal{test}}})}=\sum_{v\in\mathcal{V}(j_{u_{\textnormal{test}}})}N_{v}$ denotes the total samples from users in $u_{\text{test}}$’s cluster.
Remark 5.3 (Suboptimality Analysis).
Here $N^{\mathcal{V}(j_{u_{\textnormal{test}}})}$ depends on both the number of users in the cluster and the samples per user. Theorem 5.2 matches the lower bound discussed in Section 6. A key requirement for achieving this bound is the sufficiency of data for each user (5.1). When this assumption is violated, Off-CLUB may incorporate samples from highly biased users with preference vectors largely different with $u_{\text{test}}$, degrading performance. In contrast, as shown in Theorem 4.2, Off-C2LUB mitigates this issue by restricting heterogeneous neighbors to a much smaller set $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, controlled by the choice of $\hat{\gamma}$. Thus, while Off-CLUB performs well with sufficient data, Off-C2LUB remains a more robust algorithm in general scenarios where Assumption 5.1 does not hold. We validate this intuition with experiments in Section 7.
6 Lower Bound for Off-ClusBand and Optimality
This section presents the theoretical lower bound on suboptimality for the Off-ClusBand problem. Theorem 6.1 establishes the bound under minimal data and regularity conditions, with the proof in Appendix A.5.
@@ 6 Lower Bound for Off-ClusBand and Optimality
[…]
$\inf_{\pi}\sup_{\mathcal{Q}\in\mathcal{I}_{u_{\textnormal{test}}}}\mathbb{E}[\textnormal{SubOpt}_{\mathcal{Q}}(u_{\textnormal{test}},\pi,A_{\textnormal{test}})]\geq\sqrt{\frac{8d}{N^{\mathcal{V}(j_{u_{\textnormal{test}}})}}}.$
Remark 6.2 (Comparison under Sufficient Data).
The suboptimality bound in Theorem 5.2, $\tilde{O}\left(\sqrt{\frac{d}{N^{\mathcal{V}(j_{u_{\text{test}}})}}}\right)$, matches the lower bound in Theorem 6.1 up to logarithmic factors, proving that Algorithm 2 (Off-CLUB) achieves near-optimal performance when the dataset satisfies Assumption 5.1. Similarly, Algorithm 1 (Off-C2LUB) with $\hat{\gamma}=\gamma$ also meets the lower bound under the same assumption with the input $\alpha=1$. Consequently, all users in the same cluster as $u_{\text{test}}$ are correctly identified, guaranteeing optimal sample utilization.
Remark 6.3 (Comparison under Limited Data).
When data is insufficient, Algorithm 2 (Off-CLUB) often clusters heterogeneous users together, incorporating biased samples and degrading decision quality. In contrast, Algorithm 1 (Off-C2LUB) with $\hat{\gamma}=\gamma$ mitigates this issue by excluding biased samples, as shown in Corollary 4.5. Yet, limited data prevents it from identifying all homogeneous users correctly of the test user, leading to a minor gap between its upper bound and the theoretical lower bound in Theorem 6.1. Moreover, as shown in Theorem 4.2, when $\gamma$ is unknown, Algorithm 1 restricts the heterogeneous neighbors of $u_{\text{test}}$ to the set $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, which includes only those users whose distance to $u_{\text{test}}$ is at most $\hat{\gamma}$. In contrast, Algorithm 2 may retain connections between highly heterogeneous users due to insufficient data, increasing bias.

### v2

@@ 4.3 Strategies for Choosing $\hat{\gamma}$
As demonstrated in Theorem˜4.4, the performance of Algorithm 1 critically depends on the choice of $\hat{\gamma}$ where a carefully chosen $\hat{\gamma}$ balances noise and bias. Below we outline strategies for selecting $\hat{\gamma}$ under different scenarios. Figure 1 shows the influence of different choices for $\hat{\gamma}$.
Case 1: Exact Information of $\gamma$. We set $\hat{\gamma}=\gamma$ when the exact value of $\gamma$ is known. This ensures $\varepsilon=0$ in Theorem˜4.4, so that only users from the same true cluster as $u_{\textnormal{test}}$ are included, i.e., $\eta_{\mathcal{W}_{\gamma}(u_{\textnormal{test}})}=0$. The resulting suboptimality is formalized in Corollary 4.7.
@@
With the same notations as Theorem˜4.4, if $\hat{\gamma}=\gamma$ in Algorithm 1, the suboptimality gap satisfies:
$\textnormal{SubOpt}(u_{\textnormal{test}},\textnormal{Off-C${}^{2}$LUB},\mathcal{A}_{\textnormal{test}})\leq\tilde{O}\left(\sqrt{\frac{d}{\tilde{\lambda}_{a}N_{\mathcal{V}_{\gamma}(u_{\textnormal{test}})}}}\right).$
@@
Corollary 4.7 implies that Algorithm 1 achieves suboptimality only with noise. This is because no users from outside $u_{\textnormal{test}}$’s true cluster are included, eliminating bias entirely. Furthermore, Algorithm 1 only connects homogeneous users based on similar preference vectors, avoiding the biased inclusion of heterogeneous users. Note that choosing $\hat{\gamma}<\gamma$ can also yield a bias-free suboptimality as in Corollary˜4.7.
@@
$\textstyle\Gamma(u,v)=\left\|\hat{\bm{\theta}}_{u}-\hat{\bm{\theta}}_{v}\right\|_{2}$ $-\alpha(\text{CI}_{u}+\text{CI}_{v}),\ \tilde{\Gamma}(u,v)=\left\|\hat{\bm{\theta}}_{u}-\hat{\bm{\theta}}_{v}\right\|_{2}+\alpha(\text{CI}_{u}+\text{CI}_{v}),$ (4)
$M(u)=\{v\in\mathcal{V}\setminus\{u\}\mid\Gamma(u,v)>0\},$
where $\text{CI}_{u}$ is as defined in Equation˜1. Since $\Gamma(u,v)\leq\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}$ for $\alpha\geq 1$, it serves as a lower bound on the true preference gap. If $\Gamma(u,v)\leq 0$, $u$ and $v$ are either in the same cluster or lack sufficient data to be distinguished. Hence $M(u)$ contains users from different clusters, while those outside are either homogeneous or uncertain. When $M(u_{\text{test}})=\emptyset$, it is natural to set $\hat{\gamma}=0$, reducing Algorithm 1 to LinUCB (Abbasi-Yadkori et al., 2011) based solely on $\mathcal{D}_{u_{\text{test}}}$.
Below, we propose two policies for choosing $\hat{\gamma}$: the underestimation policy selects a smaller value to limit bias but at the cost of more noise, while the overestimation policy makes the opposite trade-off:
@@
$\hat{\gamma}=\begin{cases}\mathbb{I}\{M(u_{\textnormal{test}})\neq\emptyset\}\cdot\min_{v\in M(u_{\textnormal{test}})}\Gamma(u_{\textnormal{test}},v),&\textbf{(Underestimation Policy)},\\[6.0pt] \mathbb{I}\{M(u_{\textnormal{test}})\neq\emptyset\}\cdot\min_{v\in M(u_{\textnormal{test}})}\tilde{\Gamma}(u_{\textnormal{test}},v),&\textbf{(Overestimation Policy)},\end{cases}$ (5)
Remark 4.9 (Suitable Scenarios for Both Policies).
The key difference between the two policies lies in whether we underestimate ($\Gamma(u,v)$) or overestimate ($\tilde{\Gamma}(u,v)$) the preference gap between two users. Intuitively, the underestimation policy selects a relatively small $\hat{\gamma}$, which reduces bias but increases noise according to Lemma˜4.2, while the overestimation policy makes the opposite tradeoff. Therefore, the underestimation policy is more suitable when avoiding bias is more critical, making it particularly valuable in safety-critical applications such as medical treatment or financial decision-making, where even small biases could lead to harmful outcomes. By contrast, the overestimation policy is preferable when aggregating large, diverse data is essential. For example, in recommender systems with high-dimensional user preferences, pooling additional data helps stabilize estimates and improves accuracy, even at the cost of introducing moderate bias. Detailed theoretical analysis and further discussion of both policies are deferred to Appendix B.2 due to space constraints.
5 Algorithm for Sufficient Dataset: Off-CLUB
@@
When $\gamma$ is unknown, Off-C2LUB faces a trade-off: a quite small $\hat{\gamma}$ increases noise, while a large one introduces bias, as shown in Section 4.3. This motivates us to propose Offline Clustering of Bandits (Off-CLUB) that avoids estimating $\gamma$ and achieves strong performance under sufficient data.
@@ 5.1 Algorithm 2: Off-CLUB
Off-CLUB shares the same two-phase structure as Off-C2LUB, consisting of a Cluster Phase and a Decision Phase. Due to space constraints, full details and pseudo-code (Algorithm 2) are deferred to Appendix F. The main difference lies in graph construction: Off-CLUB starts with a complete graph and removes edges between users with dissimilar preferences, whereas Off-C2LUB begins with an empty graph and adds edges only when sufficient similarity is detected.
@@
$\textstyle\big\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\big\|_{2}>\alpha\left(\text{CI}_{u_{1}}+\text{CI}_{u_{2}}\right),$ (6)
indicating that users $u_{1}$ and $u_{2}$ belong to different clusters with high probability. After edge removal, the resulting graph $\tilde{\mathcal{G}}$ correctly clusters users with sharing preference vectors, provided that all users have sufficient samples as discussed in Section 5.2. Finally, the subsequent steps, including data aggregation, computation of statistics, and the Decision Phase, are the same to those in Off-C2LUB.
@@ 5.2 Upper Bound for Off-CLUB under Sufficient Data
The effectiveness of Off-CLUB relies on accurately removing edges between users in different clusters, requiring $\text{CI}_{u}$ to be sufficiently small. This is typically satisfied when users have many samples, while users with few samples may remain incorrectly connected, grouping heterogeneous users together. We therefore formalize the data sufficiency condition for the success of Off-CLUB:
@@
The dataset $\mathcal{D}_{u}$ is called $\delta$-sufficient for user $u$ if $N_{u}\geq\max\left\{\frac{16}{\tilde{\lambda}_{a}^{2}}\log\left(\frac{8dU}{\tilde{\lambda}_{a}^{2}\delta}\right),\,\frac{512d}{\gamma^{2}\tilde{\lambda}_{a}}\log\left(\frac{2U}{\delta}\right)\right\}$ for some $\delta\in(0,1)$. Furthermore, we assume that the full dataset $\mathcal{D}=\{\mathcal{D}_{u}\}_{u=1}^{U}$ is completely $\delta$-sufficient, i.e., each user’s dataset $\mathcal{D}_{u}$ is $\delta$-sufficient.
˜5.1 guarantees data sufficiency to estimate preferences and cluster accurately. Theorem 5.2 (proof in Appendix C.5) guarantees the performance for Off-CLUB under this assumption.
@@
For any $\delta\in(0,\frac{1}{2})$, assume $\mathcal{D}$ is completely $\delta$-sufficient. Then for $\alpha=1$, $\lambda>0$, with probability at least $1-2\delta$, the suboptimality gap satisfies
$\textnormal{SubOpt}(u_{\textnormal{test}},\textnormal{Off-CLUB},\mathcal{A}_{\textnormal{test}})\leq\tilde{O}\left(\sqrt{\frac{d}{\tilde{\lambda}_{a}N_{\mathcal{V}(j_{u_{\textnormal{test}}})}}}\right).$
Remark 5.3 (Analysis of Theorem˜5.2).
A key requirement for achieving this bound is the sufficiency of data for each user (˜5.1). When this assumption is violated, Off-CLUB may incorporate samples from highly biased users with preference vectors largely different with $u_{\text{test}}$, degrading performance. In contrast, as shown in Theorem˜4.4, Off-C2LUB mitigates this issue by restricting heterogeneous neighbors to a much smaller set $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, controlled by $\hat{\gamma}$. Thus, while Off-CLUB performs well with sufficient data, Off-C2LUB remains a more robust algorithm in general scenarios where Assumption 5.1 does not hold. We validate this with experiments in Section 6.
5.3 Lower Bound for Off-ClusBand Problem and Optimality
To facilitate further comparison between Off-CLUB’s and Off-C2LUB’s performance, we derive a theoretical lower bound for the Off-ClusBand problem. The proof is in Appendix C.6.
@@ 5.3 Lower Bound for Off-ClusBand Problem and Optimality
Given any test user $u_{\textnormal{test}}\in\mathcal{U}$, consider the set of Off-ClusBand instances: $\mathcal{I}_{u_{\textnormal{test}}}=\big\{(\mathcal{D},\Theta,R)\,\big|\,\big\|\left(M^{\mathcal{V}(j_{u_{\textnormal{test}}})}\right)^{-\frac{1}{2}}\bm{a}_{\textnormal{test}}^{*}\big\|_{2}\leq 2\sqrt{2},\;R\textnormal{ is 1-subgaussian}\big\}$, where $\Theta=\{\bm{\theta}_{u}\}_{u\in\mathcal{V}}$, $\bm{a}_{\textnormal{test}}^{*}$ is the optimal action for user $u_{\textnormal{test}}$, and $M^{\mathcal{V}(j_{u_{\textnormal{test}}})}=\sum_{v\in\mathcal{V}(j_{u_{\textnormal{test}}})}M_{v}$. If the dataset satisfies $N_{u_{\textnormal{test}}}\geq 8d$, the following lower bound holds:
$\inf_{\pi}\sup_{\mathcal{Q}\in\mathcal{I}_{u_{\textnormal{test}}}}\mathbb{E}[\textnormal{SubOpt}_{\mathcal{Q}}(u_{\textnormal{test}},\pi,\mathcal{A}_{\textnormal{test}})]\geq\sqrt{8d/N_{\mathcal{V}(j_{u_{\textnormal{test}}})}}.$
Remark 5.5 (Comparison Between Upper and Lower Bounds).
The lower bound in Theorem˜5.4 shows that the optimal strategy for Off-ClusBand is to use only samples from users in the same cluster as $u_{\text{test}}$ while excluding heterogeneous users. Recall that under Assumption 5.1, Theorem 5.2 gives a bound scaling with $1/\sqrt{N_{\mathcal{V}(j_{u_{\text{test}}})}}$. Under the same assumption, Algorithm 1 (Off-C2LUB) with $\hat{\gamma}=\gamma$ achieves the same result as Theorem 5.2 by only connecting all homogeneous users (as shown in Lemma˜4.2 with $\alpha=1$). This matches the lower bound up to logarithmic factors in terms of the total number of utilized samples $N_{\mathcal{V}(j_{u_{\text{test}}})}$, demonstrating near-optimal performance of both algorithms under Assumption 5.1. When data is sparse, however, Off-CLUB often clusters highly heterogeneous users and introduces bias, whereas Off-C2LUB reduces this risk by restricting heterogeneous neighbors to $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, explaining the strength of Off-C2LUB with insufficient data.